# Traffic Scene Generation from Natural Language Description for Autonomous Vehicles with Large Language Model

### v1

@@ IV-C Ablation Study
The primary objective of our framework design is to provide a reliable pipeline to generate traffic scenes. Therefore, we evaluate the efficacy of different components by adopting ten distinct prompts under diverse scenarios. The scenarios encompass three normal, four critical, and three specific road conditions, such as the presence of traffic lights or the absence of crossroads or a stop sign on the road.
@@ IV-C2 Scene Quality
IV-D Diversity Test
@@
We assess the diversity of these scenarios by generating each scene five times and compute the metrics, including Agent Diversity (AD) and Road Diversity (RD). These diversity metrics are calculated as the ratio of unique objects to the total number of objects. For agents, diversity considers variations in agent type, action, and relative position; for roads, diversity is assessed by unique road IDs (different directions on the same road is considered the same). Additionally, to verify the practicality and accuracy of our rendering interface, we include the metric Text Matching (TM), which evaluates whether the generated scenes accurately reflect the input prompts using a binary matched/unmatched criterion. The comprehensive results are detailed in Table II.
@@
V Conclusion and Future Work
@@
In this paper, we introduce a novel framework that leverages natural language descriptions for traffic scene generation facilitated by a Large Language Model. Our pipeline is able to automatically identify diverse and optimal locations for scene generation and allows customized options. Our pipeline starts by analyzing user input to decompose the requirements, performing road candidate retrieval, and planning detailed agents. To ensure the selected scene is suitable for an agent to act and be spawned, we adopt a road ranking strategy to match all the conditions. We provide experiments on diversity tests and adversarial training on the critical scenes to examine our pipeline ability to generate diverse and reliable scenes. Looking ahead, we plan to enhance the system’s practicality for agent and road generation to enable the creation of non-existent signals and objects and design a generated model that can transfer simulated scenes into a real-world video. This will further expand our framework’s ability to create a real and dynamic scene.

### v2

@@ IV-C Ablation Study
The primary objective of our framework design is to provide a reliable pipeline to generate traffic scenes. Therefore, we evaluate the efficacy of different components by adopting ten distinct prompts under diverse scenarios. The scenarios encompass 3 normal, four critical, and three specific road conditions, such as the presence of traffic lights or the absence of crossroads or a stop sign on the road.
@@ IV-C2 Scene Quality
V Discussion
To further clarify our approach, we conduct additional experiments to discuss the diversity (Sec. V-A) and demonstrate an extra application in driving captioning (Sec. V-B).
V-A Diversity Test
@@
We assess the diversity of these scenarios by generating each scene five times and compute the metrics, including Agent Diversity (AD) and Road Diversity (RD). These diversity metrics are calculated as the ratio of unique objects to the total number of objects. For agents, diversity considers variations in agent type, action, and relative position; for roads, diversity is assessed by unique road IDs (different directions on the same road are considered the same). Additionally, to verify the practicality and accuracy of our rendering interface, we include the metric Text Matching (TM), which evaluates whether the generated scenes accurately reflect the input prompts using a binary matched/unmatched criterion. The comprehensive results are detailed in Table II.
@@
V-B Driving Captioning
Since our pipeline provides images from the car’s front view, it can also be utilized for training driving captioning models for action reasoning under critical scenes. Captioning models, such as ADAPT [7] and BehaviorGPT4 [29], generate narrations of driving actions and their reasoning based on video input. However, as these models are trained on real-world datasets like BDD-X [41], they lack exposure to critical scenarios.
We illustrate an example in Fig. 6. In this case, the expected action is “turns left slowly,” with the reason of “giving way to a coming car.” However, the model fails to capture “slow” and only describes the left turn. Additionally, the generated reasoning is incorrect due to hallucination, further indicating that the original model does not effectively account for critical scenarios.
To improve the model’s critical reasoning capabilities, we sample 20 critical scenes for training and another 20 for evaluation from our pipeline. To ensure the collected scenes accurately reflect how the ego-agent responds to critical situations, we set its behavior to be cautious and resample whenever a collision occurs. Since the captions generated by our pipeline do not naturally follow a narration-with-reasoning structure, we use an LLM to reformat the scene descriptions to match the training format. For our experiments, we select the model from ADAPT [7], as it does not rely on closed-source models like GPT-4 and can be trained efficiently without excessive computational resources.
The results in Tab. IV are evaluated using standard captioning metrics, including BLEU [42], METEOR [6], and CIDEr [43], on our critical scenario data. Before fine-tuning, the model performs significantly worse on critical scenarios, particularly in reasoning. It struggles to recognize critical objects and fails to infer the appropriate actions. In contrast, after learning from a small set of examples, the model improves significantly in reasoning about actions, allowing it to generate more reliable descriptions for abnormal or critical scenes. Additionally, the descriptions of actions become more precise, leading to better captioning ability.
Figure 6: Illustration of predicted captions under critical scenario generation. N denotes narration, while R represents reasoning. The subject of narration is omitted as it always refers to the ego-car.
TABLE IV: Results of Driving Captioning on critical scenario.
Narration Reasoning
B M C B M C
original 4.8 13.5 15.2 0.0 10.0 18.4
finetuned 9.9 18.4 52.9 7.2 11.2 51.9
VI Failure Analysis
In this section, we analyze the failure cases observed in our experiments and discuss potential reasons and solutions.
VI-A Planning Failures
Since our framework is built upon a Large Language Model (LLM), the generated scenes highly depend on the LLM agent’s planning capabilities. Although we incorporate a reasoning stage to enhance the accuracy of both the agent and the road conditions, the LLM can still produce incorrect results in certain cases. One common issue is the LLM’s tendency to overlook alternative phrasing. For example, if the input prompt describes the scenario as “a car in front does not move for an infinite amount of time,” the LLM agent may fail to recognize this as a blocking action. Another challenge arises from hallucinations generated by the LLM agent. For instance, when instructed to generate two or more cars, the model frequently introduces a traffic light as a required signal for the road, even when it is not specified in the prompt. To mitigate these issues, one possible approach is to employ a reasoning LLM agent such as OpenAI-o1 [44] to provide better planning results. Additionally, fine-tuning an open-source LLM using expert demonstrations or preference alignment [45] can help reduce hallucinations and improve overall accuracy.
VI-B Retrieval Failures
Another source of failure arises from missing signals and objects within the CARLA environment. Since our framework relies on the built-in map, complex combinations of signals and objects and novel conditions cannot be retrieved. For instance, identifying a road that includes left, right, and straight directions while excluding traffic lights and stop signs is currently infeasible. This limitation exists because all all-directional intersections are designed to include either a traffic light or a stop sign. Additionally, while a speed limit sign displaying “70” is supported, requesting a speed limit sign of “200” is not. Similar limitations arise when users specify conditions that are not pre-defined in the database. To overcome these constraints, we plan to enhance support for generating non-existent signals, objects, and roads by directly creating OpenDRIVE code and integrating it into the map. This approach will allow users to customize various signals and objects, enabling more diverse scene generation.
VII Conclusion and Future Work
@@
In this paper, we introduce a novel framework that leverages natural language descriptions for traffic scene generation facilitated by a Large Language Model. Our pipeline is able to understand scenario requirements effectively and find the best virtual environment locations for scene generation with diverse and customized options. Our pipeline starts by analyzing user input to decompose the requirements, performing road candidate retrieval, and planning detailed agents. To ensure the selected scene is suitable for an agent to act and be spawned, we adopt a road ranking strategy to match all the conditions. We provide experiments on diversity tests and adversarial training on the critical scenes to examine our pipeline ability to generate diverse and reliable scenes. Looking ahead, we plan to enhance the system’s practicality for agent and road generation to enable the creation of non-existent signals and objects and design a generated model that can transfer simulated scenes into a real-world video. This will further expand our framework’s ability to create a real and dynamic scene.